OpenStudy (anonymous):

Evaluate b^2c^-1 for b = -4 and c = 2. is it -8

OpenStudy (anonymous):

There is minus mistake: \[(-4)^2 \times \frac{1}{2} = 16 \times \frac{1}{2} = ??\]

OpenStudy (jiteshmeghwal9):

plug the given values in the expression.:)

OpenStudy (anonymous):

oh.. 8

OpenStudy (anonymous):

c^-1=1/c so: b^2c^-1 = b^2/c = 16/2

OpenStudy (anonymous):

Yes now you are right..

OpenStudy (anonymous):

\[\large (-a)^2 = a^2\]

OpenStudy (jiteshmeghwal9):

\[\Huge{\color{red}{\checkmark}}\]@helpme9283